# To Tag, or Not to Tag: Translating C’s Unions to Rust’s Tagged Unions

### v1

@@ Abstract.
Automatic C-to-Rust translation is a promising way to enhance the reliability of legacy system software. However, C2Rust, an industrially developed translator, generates Rust code with unsafe features, undermining the translation’s objective. While researchers have proposed techniques to remove unsafe features in C2Rust-generated code, these efforts have targeted only a limited subset of unsafe features. One important unsafe feature remaining unaddressed is a union, a type consisting of multiple fields sharing the same memory storage. Programmers often place a union with a tag in a struct to record the last-written field, but they can still access wrong fields. In contrast, Rust’s tagged unions combine tags and unions at the language level, ensuring correct value access. In this work, we propose techniques to replace unions with tagged unions during C-to-Rust translation. We develop a static analysis that facilitates such replacement by identifying tag fields and the corresponding tag values. The analysis involves a must-points-to analysis computing struct field values and a heuristic interpreting these results. To enhance scalability, we adopt intraprocedural function-wise analysis, allowing selective analysis of functions. Our evaluation on 35 real-world C programs shows that the proposed approach is (1) precise, identifying 74 tag fields with no false positives and only five false negatives, (2) mostly correct, with 17 out of 23 programs passing tests post-transformation, and (3) scalable, capable of analyzing and transforming 192k LOC in 4,910 seconds.
@@ 1. Introduction
Translating C code to Rust is a promising approach to enhancing the reliability of legacy system software. Programs written in C often suffer from memory bugs leading to critical security vulnerabilities due to the absence of language-level mechanisms to prevent them (Chen et al., 2011; Thomas, 2019). Rust, a recently developed system programming language, ensures memory safety at compile time through type checking (Matsakis and Klock, 2014; Jung et al., 2017). By translating legacy C code to Rust, developers can detect previously unknown bugs and prevent introducing new bugs (Hutt, 2021).
Since manual translation is laborious and error-prone, an automatic C-to-Rust translator named C2Rust (Wingerter, 2022) has been developed in the industry. It converts C code to Rust by leveraging Unsafe Rust (Rust, 2024f), which allows the use of unsafe language features. These features, such as dereferencing raw pointers and calling functions in external code, are equivalent to C’s features and enable straightforward syntactic translation. However, as the compiler does not ensure their safety, their use contradicts the goal of translation.
@@
Unions are an important source of unsafety in C-to-Rust translation that has not been studied yet. A union is a compound data type consisting of multiple fields sharing the same memory storage, facilitating efficient memory use by allowing values of different types to be stored at the same location (Rust, 2024g). Since memory efficiency is crucial in system software, unions are widely used in C. Notably, Emre et al. (Emre et al., 2021) show that 18% of unsafe functions (functions using unsafe features) in C2Rust-generated code involve unions.
@@
The second requirement for the analysis is scalability. The key idea for achieving scalability is to selectively analyze functions. To identify tag fields, we need the field values only of the structs containing unions. It means that functions not accessing such structs do not need to be analyzed. Therefore, we adopt intraprocedural function-wise analysis, instead of interprocedural whole-program analysis, allowing only the selected functions to be analyzed.
@@
  • We implement the proposed approach in a tool named Urcrat (union-removing C-to-Rust automatic translator) and evaluate it using 35 real-world C programs. Our evaluation shows that the approach is (1) precise, identifying 74 tag fields with no false positives and only five false negatives, (2) mostly correct, with 17 out of 23 programs passing tests after transformation, and (3) scalable, capable of analyzing and transforming 192k LOC in 4,910 seconds (Section 5).
@@ 3.1. Candidate Identification
To concretely define candidate structs, we first define tag-eligible fields, those that can potentially serve as tag fields. A field of a struct is considered tag-eligible if (1) it has an integer type, (2) when it appears on the left side of an assignment, the operator is =, excluding others such as +=, and (3) it is never referenced by a pointer. The second and third conditions arise from limitations in the expressibility of tagged unions. Tags of tagged unions are not integers and thus cannot undergo integer operations. Moreover, since tags do not exist as fields, they cannot be referenced. Consequently, fields that exhibit such behavior cannot be transformed into tagged unions.
@@ 3.3. Must-Points-To Analysis
We first describe how we visualize graphs throughout this section. Since nodes denote memory locations, some nodes correspond to the stack locations used by local variables. For clarity in visualization, we draw an arrow from the name of a local variable to the node representing its memory location. Note that a variable name is not a node and thus this arrow is not an edge. For example, x = &y constructs the following graph:
@@
Using the @$N$ label is beneficial because it enables efficient propagation of integer values to memory locations known to hold the same value. Consider the following graph, constructed by x = y:
@@ 4. Code Transformation
This result is the same as the hand-written code in Section 2.3, except for the variant names. Since we cannot automatically determine suitable names for each variant, we generate names by concatenating the union field name with the tag value. For tags not associated with any fields, we prepend Empty to the tag value. Although these generated names may reduce code readability, developers can easily rename them to more meaningful names using IDEs.
@@ 4.2. Idiomatic Transformation
However, the idiomatic approach is not applicable to other kinds of conditions, such as conjunctions and inequalities, since they cannot be represented with pattern matching on tagged union values. Thus, the following code is transformed using the naïve approach:
@@ 5. Evaluation
In this section, we evaluate our approach with 35 real-world C programs. We first describe our implementation of Urcrat, which realizes the proposed approach (Section 5.1), and the process of collecting the benchmark programs (Section 5.2). We then assess our approach by addressing the following research questions:
  • RQ1. Precision: Does it identify tag fields without false positives or false negatives? (Section 5.3)
@@
  • RQ3. Scalability: Does it efficiently analyze and transform large programs? (Section 5.5)
@@ 5.2. Benchmark Program Collection
We collected benchmark programs from three sources: (1) Crown (Zhang et al., 2023), (2) GNU packages (GNU, 2024), and (3) GitHub. Only 2 out of 20 programs used by Crown have candidate unions, so we expanded the benchmarks with additional sources. We chose GNU packages for their representative C projects and GitHub for its diverse code patterns. From GNU, we gathered C packages with less than 100k LOC and individual Wikipedia entries, indicating they are well-known. From GitHub, we gathered C projects with less than 1 MB of code and over 1,000 stars. In both collections, we retained programs (1) compiled successfully on Ubuntu, (2) transpiled successfully by C2Rust, and (3) containing candidate unions. This resulted in 23 programs from GNU and 10 from GitHub, giving us a total of 35 benchmark programs. We avoided large codebases because C2Rust often produces Rust code with type errors, primarily due to missing type casts, which require significant manual effort to correct. Columns 2–5 of Table 1 present the C LOC, Rust LOC, number of unions, and number of candidate unions in each program, respectively.
5.3. RQ1: Precision
We evaluate the precision of the proposed approach. We first identified the tag field of each candidate union through static analysis. Column 6 of Table 1 shows the number of candidate unions for which a tag field is identified in each benchmark program. Out of 127 candidates across 35 programs, Urcrat identifies tag fields for 74 candidates in 29 programs. We then manually inspected each candidate union to determine the presence of a tag field, checking for false positives (a tag field identified for a union that does not have one) and false negatives (no tag field identified for a union that has one). For practical use of our tool, false positives are problematic because they change the program’s semantics. Conversely, false negatives are less problematic as they only prevent the replacement of unions with tagged unions, and the semantics remains correct.
Our manual inspection shows that the static analysis is precise, revealing only five false negatives and no false positives. Specifically, two false negatives occur in glpk-5.0, while the others are in gawk-5.2.2, screen-4.9.0, and uucp-1.07, respectively. Three of these (from glpk-5.0, gawk-5.2.2, and screen-4.9.0) are due to intermediate states involving tag values that are not associated with any union fields by CollectFromAccesses. We now discuss the reasons for the remaining two false negatives:
@@ 5.4. RQ2: Correctness
5.5. RQ3: Scalability
@@
We evaluate the scalability of the proposed approach by measuring the execution time of Urcrat, which includes both analysis and transformation, for each program. Fig. 2 presents the execution time relative to the Rust LOC, with the y-axis displayed in a log scale due to the wide range of execution times. Urcrat efficiently handles most programs, with 31 programs taking less than a minute. The longest execution time is 4,910 seconds for gawk-5.2.2. Execution time shows only a weak correlation with code size, as other factors, primarily the complexity of pointer graphs and the number of analyzed functions, also significantly influence execution time.
We also investigate the effectiveness of our selective function analysis in terms of scalability. For each program, the transformation consumes less than 1% of the total time, as it involves only tree-walking, while the analysis occupies the remaining time. This highlights the importance of reducing analysis time. We calculated the percentage of analyzed functions relative to the total functions in each program, finding a geometric mean of 4.77%. In 27 programs, the analyzed functions constitute less than 10% of the total functions. The highest percentage is 58.23% in the case of minilisp. These results indicate that our approach is effective for scalability.
@@ 5.7. RQ5: Impact on Performance
The experimental results show that the performance overhead is negligible. On average, the transformed programs were only 0.01% slower than the original ones. Among the 20 programs, 5 ran slightly slower after transformation, while the others were faster.
@@ 5.8. Threats to Validity
Threats to external validity stem from the choice of benchmarks. GNU packages often share common code patterns, which may introduce bias. Although we included projects from GitHub to enhance diversity, this may not fully represent the entire C ecosystem. Notably, large codebases might exhibit different characteristics. Further experiments with a broader range of C programs would provide greater confidence in the generalizability of our approach.

### v2

@@ Abstract.
Automatic C-to-Rust translation is a promising way to enhance the reliability of legacy system software. However, C2Rust, an industrially developed translator, generates Rust code with unsafe features, undermining the translation’s objective. While researchers have proposed techniques to remove unsafe features in C2Rust-generated code, these efforts have targeted only a limited subset of unsafe features. One important unsafe feature remaining unaddressed is a union, a type consisting of multiple fields sharing the same memory storage. Programmers often place a union with a tag in a struct to record the last-written field, but they can still access wrong fields. In contrast, Rust’s tagged unions combine tags and unions at the language level, ensuring correct value access. In this work, we propose techniques to replace unions with tagged unions during C-to-Rust translation. We develop a static analysis that facilitates such replacement by identifying tag fields and the corresponding tag values. The analysis involves a must-points-to analysis computing struct field values and a heuristic interpreting these results. To enhance efficiency, we adopt intraprocedural function-wise analysis, allowing selective analysis of functions. Our evaluation on 36 real-world C programs shows that the proposed approach is (1) precise, identifying 74 tag fields with no false positives and only five false negatives, (2) mostly correct, with 17 out of 23 programs passing tests post-transformation, and (3) efficient, capable of analyzing and transforming 141k LOC in 4,910 seconds.
Rust, C, Automatic Translation, Union, Tagged Union
journalyear: 2024copyright: rightsretainedconference: 39th IEEE/ACM International Conference on Automated Software Engineering ; October 27-November 1, 2024; Sacramento, CA, USAbooktitle: 39th IEEE/ACM International Conference on Automated Software Engineering (ASE ’24), October 27-November 1, 2024, Sacramento, CA, USAdoi: 10.1145/3691620.3694985isbn: 979-8-4007-1248-7/24/10ccs: Software and its engineering Source code generationccs: Software and its engineering Automated static analysisccs: Software and its engineering Maintaining softwareccs: Software and its engineering Software evolution
@@ 1. Introduction
Translating C code to Rust is a promising approach to enhancing the reliability of legacy system software. C Programs often suffer from memory bugs leading to critical security vulnerabilities due to the absence of language-level mechanisms to prevent them (Chen et al., 2011; Thomas, 2019). Rust, a recently developed system programming language, ensures memory safety at compile time through type checking (Matsakis and Klock, 2014; Jung et al., 2017). By translating legacy C code to Rust, developers can detect previously unknown bugs and prevent introducing new bugs (Hutt, 2021).
Since manual translation is laborious and error-prone, an automatic C-to-Rust translator named C2Rust (Wingerter, 2022) has been developed in the industry. It converts C code to Rust by leveraging Unsafe Rust (Rust, 2024g), which allows the use of unsafe language features. These features, such as dereferencing raw pointers and calling functions in external code, are equivalent to C’s features and enable straightforward syntactic translation. However, as the compiler does not ensure their safety, their use contradicts the goal of translation.
@@
Unions are an important source of unsafety in C-to-Rust translation that has not been studied yet. A union is a compound data type consisting of multiple fields sharing the same memory storage, facilitating efficient memory use by allowing values of different types to be stored at the same location (Rust, 2024h). Since memory efficiency is crucial in system software, unions are widely used in C. Notably, Emre et al. (Emre et al., 2021) show that 18% of unsafe functions (functions using unsafe features) in C2Rust-generated code involve unions.
@@
The second requirement for the analysis is efficiency. The key idea for achieving efficiency is to selectively analyze functions. To identify tag fields, we need the field values only of the structs containing unions. It means that functions not accessing such structs do not need to be analyzed. Therefore, we adopt intraprocedural function-wise analysis, instead of interprocedural whole-program analysis, allowing only the selected functions to be analyzed.
@@
  • We implement the proposed approach in a tool named Urcrat (union-removing C-to-Rust automatic translator) and evaluate it using 36 real-world C programs. Our evaluation shows that the approach is (1) precise, identifying 74 tag fields with no false positives and only five false negatives, (2) mostly correct, with 17 out of 23 programs passing tests after transformation, and (3) efficient, capable of analyzing and transforming 141k LOC in 4,910 seconds (Section 5).
@@ 3.1. Candidate Identification
To concretely define candidate structs, we first define tag-eligible fields, those that can potentially serve as tag fields. A field of a struct is considered tag-eligible if (1) it has an integer type, (2) when it appears on the left side of an assignment, the operator is =, excluding others such as +=, and (3) it is never referenced by a pointer. For the first condition, integer types include bool, i8, u8, i16, u16, i32, u32, i64, and u64, which C2Rust translates from _Bool, signed/unsigned char, short, int, and long. The second and third conditions arise from limitations in the expressibility of tagged unions. Tags of tagged unions are not integers and thus cannot undergo integer operations. Moreover, since tags do not exist as fields, they cannot be referenced. Consequently, fields that exhibit such behavior cannot be transformed into tagged unions.
@@ 3.3. Must-Points-To Analysis
We first describe how we visualize graphs throughout this section. Since nodes denote memory locations, some nodes correspond to the stack locations used by local variables. For clarity in visualization, we draw a dashed arrow from the name of a local variable to the node representing its memory location. Note that a variable name is not a node and thus this arrow is not an edge. For example, x = &y constructs the following graph:
@@
Using the @$N$ label is beneficial since it enables efficient propagation of integer values to memory locations known to hold the same value. Consider the following graph, constructed by x = y:
@@ 4. Code Transformation
struct BExpr { l: *mut Expr, r: *mut Expr }
This result is the same as the hand-written code in Section 2.3, except for the variant names. We generate variant names by concatenating the union field name with the tag value. For tags not associated with any fields, we prepend Empty to the tag value. To improve variant names, one option is to utilize global variables’ names. When using unions with tag fields, C programmers often define an enum, i.e., a group of constant integers with associated names, to use them as tag values. These names typically reflect the programmers’ understanding of the tag values’ meaning, e.g., EXPR_ONE. Since C2Rust translates each enum definition into multiple constant global variable definitions while preserving the names, we can use these variable names instead of the union field names. Although the names generated with this strategy may still be unsatisfactory, developers can easily rename them to more meaningful names using IDEs.
@@ 4.2. Idiomatic Transformation
However, the idiomatic approach is not applicable to other kinds of conditions, such as conjunctions and inequalities, because Rust’s patterns currently cannot represent these conditions. A conjunction specifies that the tag is a particular integer and also that some boolean formula is satisfied. When using match for pattern matching, Rust offers match guards (Rust, 2024f) to express such logic. However, when using if-let, this logic is supported through let chains (Farrokhzad, 2018), which is currently an unstable feature and requires a special flag to enable. For this reason, we chose not to transform conjunctions into pattern matching in this work, leaving it for future work once this feature is stabilized. On the other hand, an inequality specifies that the tag is not a particular integer. However, the purpose of pattern matching is to check whether a value conforms to a certain pattern, not that it does not, and thus it cannot replace inequalities. Thus, the following code is transformed using the naïve approach:
@@
The idiomatic transformation is also not applicable when union fields are accessed without using match or if to check the tag field. C programmers sometimes assume that a tag field has a specific value at a certain point based on their understanding of the program’s behavior and directly access the union field without checking the tag. Since tag check is not performed, we cannot transform such code into pattern matching and must resort to the naïve transformation.
@@ 5. Evaluation
In this section, we evaluate our approach with 36 real-world C programs. We first describe our implementation of Urcrat, which realizes the proposed approach (Section 5.1), and the process of collecting the benchmark programs (Section 5.2). The implementation and benchmark programs are publicly available (Hong and Ryu, 2024). We then assess our approach by addressing the following research questions:
  • RQ1. Precision and recall: Does it identify tag fields without false positives or false negatives? (Section 5.3)
@@
  • RQ3. Efficiency: Does it efficiently analyze and transform programs? (Section 5.5)
@@ 5.2. Benchmark Program Collection
We collected benchmark programs from three sources: (1) Crown (Zhang et al., 2023), (2) GNU packages (GNU, 2024), and (3) GitHub. Only 2 out of 20 programs used by Crown have candidate unions, so we expanded the benchmarks with additional sources. We chose GNU packages for their representative C projects and GitHub for its diverse code patterns. We avoided large codebases because C2Rust often produces Rust code with type errors, primarily due to missing type casts, which require significant manual effort to correct. From GNU, we gathered C packages with less than 250k LOC and individual Wikipedia entries, indicating they are well-known. From GitHub, we gathered C projects with less than 1 MB of code and over 1,000 stars. In both collections, we retained programs (1) compiled successfully on Ubuntu, (2) transpiled successfully by C2Rust, and (3) containing candidate unions. This resulted in 24 programs from GNU and 10 from GitHub, giving us a total of 36 benchmark programs. Of these, 20 required manual fixes after C2Rust’s translation, with an average of 26.6 lines modified. Columns 2–5 of Table 1 present the C LOC, Rust LOC, number of unions, and number of candidate unions in each program, respectively. In the benchmarks, 21% of unsafe functions involve unions, and 11% involve tagged unions.
5.3. RQ1: Precision and recall
We evaluate the precision and recall of the proposed approach. We first identified the tag field of each candidate union through static analysis. Column 6 of Table 1 shows the number of candidate unions for which a tag field is identified in each benchmark program. Out of 141 candidates across 36 programs, Urcrat identifies tag fields for 74 candidates in 29 programs. We then manually inspected each candidate union to determine the presence of a tag field, checking for false positives (a tag field identified for a union that does not have one) and false negatives (no tag field identified for a union that has one). For practical use of our tool, false positives are problematic because they change the program’s semantics. Conversely, false negatives are less problematic as they only prevent the replacement of unions with tagged unions, and the semantics remains correct.
Our manual inspection shows that the static analysis is precise, revealing no false positives and only five false negatives. This results in a precision of 100% and a recall of 93.7% ($=74/79$). Specifically, two false negatives occur in glpk-5.0, while the others are in gawk-5.2.2, screen-4.9.0, and uucp-1.07, respectively. Three of these (from glpk-5.0, gawk-5.2.2, and screen-4.9.0) are due to intermediate states involving tag values that are not associated with any union fields by CollectFromAccesses. We now discuss the reasons for the remaining two false negatives:
@@ 5.4. RQ2: Correctness
Currently, for end users of the tool, detecting and fixing incorrect translations is challenging. To identify incorrect translations, they must manually inspect the code or run tests. If test suites do not exist, creating new test cases can be costly. Additionally, test cases may not always reveal incorrect translations. Fixing incorrect code is even more challenging, as users need to identify the root cause. The difficulty of this process depends on the root cause. When tag values are not correctly identified or fields not lastly written are read, investigating the cause is relatively straightforward. The test fails due to a panic, allowing users to check which tag value or field read triggered the panic. However, issues related to different memory layouts are more difficult to diagnose since they do not trigger panics. In such cases, users must rely on their debugging skills. We believe it would be beneficial to identify common incorrect translation patterns and design analyses to detect these patterns, thereby providing users with warnings. We leave this as future work.
5.5. RQ3: Efficiency
@@
We evaluate the efficiency of the proposed approach by measuring the execution time of Urcrat, which includes both analysis and transformation, for each program. Fig. 2 presents the execution time relative to the Rust LOC, with the y-axis displayed in a log scale due to the wide range of execution times. Urcrat efficiently handles most programs, with 31 programs taking less than a minute. The longest execution time is 4,910 seconds for gawk-5.2.2. Execution time shows only a weak correlation with code size, as other factors, primarily the complexity of pointer graphs and the number of analyzed functions, also significantly influence execution time.
We also investigate the effectiveness of our selective function analysis in terms of efficiency. For each program, the transformation consumes less than 1% of the total time, as it involves only tree-walking, while the analysis occupies the remaining time. This highlights the importance of reducing analysis time. We calculated the percentage of analyzed functions relative to the total functions in each program, finding a geometric mean of 4.80%. In 28 programs, the analyzed functions constitute less than 10% of the total functions. The highest percentage is 58.23% in the case of minilisp. These results indicate that our approach is effective for efficiency.
@@ 5.7. RQ5: Impact on Performance
The experimental results show that the performance overhead is negligible. On average, the transformed programs were only 0.01% slower than the original ones. Among the 20 programs, 5 ran slightly slower after transformation, while the others were faster. We also performed one-sided t-test with the null hypothesis that the program after transformation is slower than the original program by at least 1%. With a significance level of 0.05, we could reject the null hypothesis for 9 out of 20 programs, implying that the transformation is unlikely to incur observable performance overhead for them. For the remaining 11 programs, we could not reject the null hypothesis, but this does not necessarily mean they have an overhead greater than 1%. Repeating the experiments may provide stronger statistical significance, rejecting the null hypothesis.
@@ 5.8. Threats to Validity
Threats to external validity primarily stem from the choice of benchmarks. GNU packages often share common code patterns, which may introduce bias. Although we included projects from GitHub to enhance diversity, this may not fully represent the entire C ecosystem. Notably, large codebases might exhibit different characteristics. Further experiments with a broader range of C programs would provide greater confidence in the generalizability of our approach.
The type errors produced by C2Rust also pose a threat to external validity. Before using our tool, all type errors in C2Rust-generated code must be resolved. Larger programs tend to exhibit more type errors after C2Rust’s translation, requiring significant effort to correct. This could discourage users from adopting both C2Rust and Urcrat, potentially impeding the widespread adoption of our approach. To mitigate this issue, we can improve C2Rust or develop techniques for fixing type errors in C2Rust-generated code.
@@ 6. Related Work
C-to-Rust Translation with Language Models
Recent studies have developed language models capable of translating code (Roziere et al., 2020; Lachaux et al., 2024; Rozière et al., 2022; Szafraniec et al., 2023; Liu et al., 2023; Wang et al., 2021; Chen et al., 2021; Feng et al., 2020; Guo et al., 2021; Touvron et al., 2023) or proposed techniques to utilize pre-trained language models for code translation (Yang et al., 2024; Pan et al., 2024). Most focus on translating between languages like C++, Java, and Python, rather than C to Rust. A notable exception is Rozière et al. (2022), which translates C++ to Rust, but their evaluation uses only small competitive programming solutions that do not involve unions, with only 21% translated correctly. Similarly, Pan et al. (2024) use GPT-4 for C-to-Rust translation, applying it to CodeNet (Puri et al., 2021), which consists of programming problem solutions, achieving 61% correctness. These results indicate that language models frequently generate code that is uncompilable or semantically incorrect, while our approach consistently produces compilable and mostly correct code.
@@ 7. Conclusion
Acknowledgements.
This research was supported by the National Research Foundation of Korea (NRF) (2022R1A2C200366011 and 2021R1A5A1021944), the Institute for Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (2022-0-00460, 2023-2020-0-01819, and 2024-00337703), and Samsung Electronics Co., Ltd (G01210570).